OpenStudy (anonymous):

A carpenter can build a toolshed three times as fast as his apprentice. Working together, they can build the shed in 6 hours. How long would it take each of them working alone to build the shed?

OpenStudy (kropot72):

Let the time taken by the carpenter to build the tool shed, working alone, be x hours. Then the apprentice, working alone, will take 3x hours. Working alone, the carpenter will complete 1/x of the tool shed in one hour and the apprentice will complete 1/(3x) of the tool shed in one hour. Working together they will complete 1/6 of the tool shed in one one hour. Therefore we can write the equation: \[\large \frac{1}{x}+\frac{1}{3x}=\frac{1}{6}\ .........(1)\] Simplifying equation (1) gives: \[\large 3x=24\] which gives x = 8. Therefore it takes the carpenter 8 hours working alone and the apprentice 24 hours working alone.